A ball is thrown off of a 44 m cliff and lands 17 away. How long was the ball in the air?



Use 9.8 m/s2, round answer to the hundredth place, and DO NOT include units.
1
Expert's answer
2019-12-16T11:12:49-0500

The vertical distance traveled by a ball


"h=\frac{gt^2}{2}"

Hence, the time of motion of the ball


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\times 44}{9.8}}=3.00\:\rm s"